# Efficient Long-duration Talking Video Synthesis with Linear Diffusion Transformer under Multimodal Guidance

### v2

@@ B.5. Impact of Audio-guidance on Diversity
To validate the impact of different fusion schemes on diversity, we serve LetsTalk as an unconditional video generator [^1] Filling audio embeddings with 00 for fusion. and compare the diversity of character actions in the generated videos. The results in Fig. 7 demonstrate that the deeper the fusion compactness, the more static the generated video. The tightest Symbiotic Fusion generates a portrait with almost no motions, Siamese Fusion produces slight motions, while the least tight Direct Fusion we adopt results in natural motions.
Appendix C Additional Implementation Details
C.1. Diffusion
We leverage a pre-trained variational autoencoder model from Stable Diffusion 1.4, with all our diffusion models operating in the latent space of this VAE. After sampling a latent from the diffusion model, we decode it to pixel space using the VAE decoder. We adopt the hyperparameters from ADM (Dhariwal and Nichol, 2021), specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$. Additionally, we use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their approach for embedding input timesteps.
C.2. Training
We initialize LetsTalk from a pre-trained Latte (Ma et al., 2024) model on the FaceForensics dataset (Rössler et al., 2018). Following the training strategy outlined in Latte, we adopt gradient clipping to stabilize training and prevent gradient explosion. For training, we use the AdamW optimizer with a constant learning rate of $1\times 10^{-4}$ and no weight decay. The batch size ranges from 10 to 16, and the only data augmentation applied is horizontal flips. Each model is trained for 300k to 500k iterations. Following common practices in generative modeling works (Peebles and Xie, 2023; Bao et al., 2023), an exponential moving average (EMA) weight with a decay rate of $0.9999$ is maintained throughout training, and all reported results are directly derived from the EMA model. Additionally, we utilize a pre-trained VAE from Stable Diffusion 1.4, which remains frozen during training. Training is conducted on both RTX 3090 and A800 GPUs, using automatic mixed precision to optimize both performance and memory usage.
C.3. Inference
During the inference stage, the network takes a single reference portrait and driving audio as input, generating a video sequence that animates the reference image based on the corresponding audio. The length and resolution of the generated video clip during inference need to remain consistent with those used during training, which in our experiments are $16$ frames and a resolution of $256\times 256$. We use DDPM for the sampling strategy, with a total of $250$ sampling steps.
C.4. Diffusion
We leverage a pre-trained variational autoencoder model from Stable Diffusion 1.4, with all our diffusion models operating in the latent space of this VAE. After sampling a latent from the diffusion model, we decode it to pixel space using the VAE decoder. We adopt the hyperparameters from ADM (Dhariwal and Nichol, 2021), specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$. Additionally, we use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their approach for embedding input timesteps.
Appendix D Qualitative Comparisons
In this section, we report comprehensive qualitative comparisons of LetsTalk with other cutting-edge methods and demonstrate more applications of LetsTalk. We show more results on the HDTF (Zhang et al., 2021) and Celeb-HQ (Zhu et al., 2022) benchmarks ( Fig. 8 and Fig. 9).
We show the portrait animation results of LetsTalk driven by English (Fig. 10), Chinese (Fig. 11), and singing audios (Fig. 12). We further use the powerful AI assistant (Doubao) to generate portraits for comparisons (Fig. 13). Note that LetsTalk is only trained on the HDTF and Celeb-HQ datasets, which only contain daily conversations of real persons.
Figure 8. The qualitative comparisons with other cutting-edge methods on the HDTF dataset.
Figure 9. The qualitative comparisons with other cutting-edge methods on the Celeb-HQ dataset.
Appendix E Visualization
We provide video visualizations in our attachment. Notably, the model used to generate the video is only trained on HDTF (Zhang et al., 2021) and CelebV-HQ (Zhu et al., 2022), which only include English conversation scenarios of real portraits. In addition, we use Chinese audio and singing audio to drive real portraits. Furthermore, LetsTalk can drive AI-generated portraits [^2] We use Doubao to generate AI-generated portraits in the attached video.. The visualization results demonstrate that LetsTalk has great generalization and achieves remarkable results outside the dataset.
Figure 10. English audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 11. Chinese audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
@@ Appendix E Visualization
Figure 12. Singing audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 13. Results of LetsTalk driving AI-generated portraits using English audio. Note that each column corresponds to the same audio.

### v3

@@ IV-F Ablation Study
Impact of Audio-guidance on Diversity To validate the impact of different fusion schemes on diversity, we serve LetsTalk as an unconditional video generator [^1] Filling audio embeddings with 0 for fusion. and compare the diversity of character actions in the generated videos. The results in Fig. 9 demonstrate that the deeper the fusion compactness, the more static the generated video. The tightest Symbiotic Fusion generates a portrait with almost no motions, Siamese Fusion produces slight motions, while the least tight Direct Fusion we adopt results in natural motions.
IV-G User Study
To comprehensively assess the perceptual quality of our approach, we conducted a user study with 12 participants. Each participant was shown videos generated by three different methods: Hallo, EchoMimic, and our proposed LetsTalk-L. For the evaluation, 20 audio samples were randomly selected, and for each audio, participants were asked to rank the three corresponding videos according to two criteria: realism, which measures how natural and lifelike the video appears, and synchronization, which evaluates how well the lip movements match the audio.
The results of the user study are summarized in Fig 10. The table reports the percentage of times each method was ranked first, second, or third for both realism and synchronization. As shown in the results, LetsTalk-L achieves the highest proportion of first-place rankings in both categories. These findings indicate that users consistently preferred the videos generated by our method, both in terms of visual realism and the accuracy of lip synchronization with the audio. Overall, the user study demonstrates the effectiveness of our approach in producing high-quality, realistic, and well-synchronized talking head videos.
Figure 10: User study results (%) on (a) realism (left) and (b) synchronization (right).
@@ IV-G User Study
Figure 11: Singing audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
IV-H Long-Duration Generation
To validate the effectiveness of our proposed approach for long video generation, we conducted experiments to analyze the quality of 100 consecutive video clips (about 64 second) generated under three settings: (1) the baseline model (LetsTalk-L) with with the memory bank, (2) the method augmented with regularized training strategy, and (3) the method further enhanced with the patch drop augmentation. Our results in Table VII demonstrate that the memory bank significantly improves temporal consistency, while the regularized training strategy further mitigates degradation in later clips, leading to higher overall video quality. The results demonstrate that regularized training strategy primarily enhances single-frame quality, significantly reducing FID from 52.243 to 32.746, while maintaining temporal consistency (FVD 264.709). On the other hand, patch drop augmentation further improves long-term motion coherence, achieving the best FVD (251.684) with only a slight FID increase (34.821), confirming its complementary role in stabilizing long-duration generation.
[…]
52.243 276.16
43.728 263.173
32.746 264.709
34.821 251.684
TABLE VII: Experimental results of long video generation. Experimental results comparing three approaches for 64-second video generation (100 clips): (1) memory bank baseline, (2) regularized training, and (3) patch drop augmentation. ✓ indicates the component is enabled.
IV-I Visualization
We provide video visualizations showcasing long-duration results of up to 2 minutes. In these videos, each row corresponds to a different portrait driven by the same audio segment, demonstrating our model’s ability to generate consistent and diverse talking head videos across various identities. Notably, our model is trained only on HDTF [55] and CelebV-HQ [56], which feature English conversational scenarios with real portraits. To further demonstrate the generalization ability of our approach, we also present results where real portraits are driven by Chinese speech and singing audio, even though such scenarios are not included in the training data.
V Conclusion
In this work, we present LetsTalk, a novel diffusion transformer framework for efficient and high-quality long-duration talking video synthesis under multimodal guidance. LetsTalk introduces a memory bank mechanism and noise-regularized training to explicitly maintain contextual continuity and mitigate error accumulation in extended video generation. Our spatiotemporal modeling, based on a deep compression autoencoder and linear attention, achieves state-of-the-art visual fidelity with substantially fewer parameters and faster inference. Through systematic analysis of multimodal fusion, we provide principled guidance for multimodal conditioning in video generation, demonstrating that deep Symbiotic Fusion for portrait features and shallow Direct Fusion for audio ensures both strong identity preservation and expressive, synchronized motion. Extensive experiments show that LetsTalk consistently outperforms existing approaches in quality, temporal coherence, and efficiency, while generalizing robustly to out-of-domain scenarios. Overall, LetsTalk advances the frontier of realistic, scalable talking video generation, paving the way for broader applications in digital humans and interactive media.